OpenStudy (anonymous):

12,11,10,9 what's next 3 terms in number sequence

OpenStudy (midhun.madhu1987):

12-1 = 11 11-1 = 10 10-1 = 9 9-1 = 8 8-1 = 7 7-1 = 6

geerky42 (geerky42):

I am surprised you cannot figure out the pattern...?

OpenStudy (midhun.madhu1987):

:P

OpenStudy (anonymous):

Oh k I get it now.....lol